OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball.

OpenStudy (amistre64):

if we assume a starting height of 0 that gives us: -16t^2 +32t; which factors to: -16t(t+2) when t = 0 or t = 2; we are back on the ground which means it would take half that time to reach its max height; when t=1 right?

OpenStudy (amistre64):

the max height whould be; but dbl chk; 42-16 = 26

OpenStudy (amistre64):

derive it to be: -32t + 32 = 0 when t=1

OpenStudy (anonymous):

a) .6 seconds to 1.8 seconds b) .2 seconds to 1.3 seconds c) .4 seconds to 1.9 seconds d) .5 seconds to 2.1 seconds

OpenStudy (amistre64):

those arent heights, those are times :)

OpenStudy (anonymous):

lol sorry i mean a) a) 10 feet b) 26 feet c) 2.3 feet d) 23 feet

OpenStudy (amistre64):

since b = 26; id go with b :)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

You need to make the variation table of the equation, 1st, the equation is defined on the interval [0, +inf[ 2nd, the derivative is h'(t)= 32 ( -t + 1 ) 3rd, its sign : + ....0.....- ( where h'(1)=0) 4th, the variation of h 5th, Hmax= h(1)=26